# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v1

@@ 6 Summary and conclusions
The parameters of the hypergeometric function and the coefficients of the two polynomials are all expressed in terms of a characteristic function $a(\ell)$, that plays the role of the SW period in a gauge theory description or the "renormalized angular momentum" in the MST formulation. The connection formulae are derived from the standard hypergeometric transformation rules. The complexity of the Heun transformation rule is hidden in the non trivial dependence $a(\ell)$ (55) on the frequency and the angular momentum $\ell$.
@@
$\begin{array}[]{lll}{\rm Near~{}horizon}&{\rm Near~{}zone}&{\rm Far~{}zone}\\ z\approx 1&x\ll z\ll 1&z\ll x\ll 1\\ F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}&G^{1}_{\alpha}=g_{\alpha}G^{0}_{\alpha}&g_{\alpha}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}\end{array}$ (106)
@@
The incoming and upgoing waveforms are given by linear combinations of any of these bases, see for example (67). When compared against the PN expansions that one can find in literature Blanchet:2013haa ; Fujita:2012cm , these formulae resum the infinite tower of logarithmically divergent terms (tail and tail of tails), responsible for the slow convergence of the PN series, into exponentials. They also summarise in a compact form the dependence on $\ell$ of the waveforms, all the so called tail and tail of tail terms of the multipole expansion and have no logarithmic terms (responsible for the slow convergence of the PN series). The results for the waveform have been compared against those obtained via MST and numerical methods, finding an agreement from 9 to 13 digits in the whole range of parameters allowed for bounded orbits.
Similar results are obtained for solutions of the non-confluent Heun equation. We remark that the results of this paper apply to any physical problem governed by a Heun differential equation or a confluence of it. In particular, they can be easily adapted to the case of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes, cosmological perturbations Bianchi:2024mlq . It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^6] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to come back to this issue in a future publication.
@@ Appendix A Gauge/Gravity/CFT dictionary
with [^7] With respect to Consoli:2022eey ; Fucito:2023afe we find it convenient to redefine the coupling sending $x\to-x$ in $Q_{CFT}$.
@@
$x=4{\rm i}\omega M\qquad,\qquad u\to(\ell+{\textstyle\frac{1}{2}})^{2}-4M^{2}\omega^{2}+10{\rm i}\omega M$ (110)

### v2

@@ VI Summary and conclusions
The parameters of the hypergeometric function and the coefficients of the two polynomials are all expressed in terms of a characteristic function $a(\ell)$, that plays the role of the SW period in a gauge theory description or the ”renormalized angular momentum” in the MST formulation. The connection formulae are derived from the standard hypergeometric transformation rules. The complexity of the Heun transformation rule is hidden in the non trivial dependence $a(\ell)$ (57) on the frequency and the angular momentum $\ell$.
@@
$\begin{array}[]{lll}{\rm Near\leavevmode\nobreak\ horizon}&{\rm Near\leavevmode\nobreak\ zone}&{\rm Far\leavevmode\nobreak\ zone}\\ z\approx 1&x\ll z\ll 1&z\ll x\ll 1\\ F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}&G^{1}_{\alpha}=g_{\alpha}G^{0}_{\alpha}&g_{\alpha}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}\end{array}$ (107)
@@
The incoming and upgoing waveforms are given by linear combinations of any of these bases, see for example (69). When compared against the PN expansions that one can find in literature [7, 20], these formulae resum the infinite tower of logarithmically divergent terms (tail and tail of tails), responsible for the slow convergence of the PN series, into exponentials. They also summarise in a compact form the dependence on $\ell$ of the waveforms, all the so called tail and tail of tail terms of the multipole expansion and have no logarithmic terms (responsible for the slow convergence of the PN series). The results for the waveform have been compared against those obtained via MST and numerical methods, finding an agreement from 9 to 13 digits in the whole range of parameters allowed for bounded orbits.
Similar results are obtained for solutions of the non-confluent Heun equation. We remark that the results of this paper apply to any physical problem governed by a Heun differential equation or a confluence of it. In particular, they can be easily adapted to the case of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes, cosmological perturbations [58]. It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^5] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to come back to this issue in a future publication.
@@ Appendix A Gauge/Gravity/CFT dictionary
with [^6] With respect to [30, 34] we find it convenient to redefine the coupling sending $x\to-x$ in $Q_{CFT}$.
@@
$x=4{\rm i}\omega M\,,\,u\to(\ell+{\textstyle\frac{1}{2}})^{2}-4M^{2}\omega^{2}+10{\rm i}\omega M$ (111)